# How Does Knowledge Selection Help Retrieval Augmented Generation?

### v3

@@ Abstract
Retrieval-augmented generation (RAG) is a powerful method for enhancing natural language generation by integrating external knowledge into a model’s output. While prior work has demonstrated the importance of improving knowledge retrieval for boosting generation quality, the role of knowledge selection remains less clear. This paper empirically analyzes how knowledge selection influences downstream generation performance in RAG systems. By simulating different retrieval and selection conditions through a controlled mixture of gold and distractor knowledge, we assess the impact of these factors on generation outcomes. Our findings indicate that the downstream generator model’s capability, as well as the complexity of the task and dataset, significantly influence the impact of knowledge selection on the overall RAG system performance. In typical scenarios, improving the knowledge recall score is key to enhancing generation outcomes, with the knowledge selector providing limited benefit when a strong generator model is used on clear, well-defined tasks. For weaker generator models or more ambiguous tasks and datasets, the knowledge F1 score becomes a critical factor, and the knowledge selector plays a more prominent role in improving overall performance.
@@ 1 Introduction
Retrieval-augmented generation (RAG) has become a pivotal technique in natural language generation, enhancing a language model’s ability to produce relevant, informed output by incorporating external knowledge Gao et al. (2023); Fan et al. (2024). This approach complements the model’s internal knowledge, which is inherently constrained by the information that was available during training, by retrieving up-to-date, relevant information to support and inform its output.
The quality of the retrieved knowledge directly influences the quality of RAG outputs. Previous studies have consistently shown that improving knowledge retrieval leads to a direct enhancement in generation performance Dinan et al. (2019); Li et al. (2022, 2024); Cuconasu et al. (2024); Wang et al. (2024); Wu et al. (2024). Similarly, effective knowledge selection has been observed to improve generation quality by filtering out retrieved information that is less relevant to the generation target Kim et al. (2020); Thulke et al. (2021); Li et al. (2022); Sun et al. (2023); Zhang et al. (2023); Zheng et al. (2024); Wang et al. (2024). However, we observe that knowledge selection is rarely used for tasks other than dialogue generation, and there are notably fewer LLM-based RAG works that use knowledge selection, compared to fine-tuned RAG models (see Section 2).
@@
Therefore, in this work, we perform a systematic empirical analysis of how knowledge selectors with various performances affect downstream RAG performance. By blending gold knowledge with distractor knowledge in varying ratios, we simulate different selection outcomes and examine their impact on the overall performance of RAG. We find that the generator model’s capability, as well as the complexity of the task and dataset, significantly influence RAG system performance. In typical scenarios, improving knowledge recall via the knowledge retriever is key to enhancing generation outcomes, with the knowledge selector providing limited additional benefit when a strong generator model is used on clear, well-defined tasks. For weaker generator models or more ambiguous tasks and datasets, knowledge F1 becomes a critical factor, and the knowledge selector plays a more prominent role in improving overall performance.
@@ 2 Related Work
has been extensively studied in recent years. Early works Guu et al. (2020); Lewis et al. (2020b); Shuster et al. (2021) jointly fine-tuned a dense retriever (e.g. DPR Karpukhin et al. (2020)) and generator (e.g. BART Lewis et al. (2020a)), which required dedicated training datasets. More recently, the introduction of large language models (LLMs) has made RAG more convenient to implement due to their strong generation performance, in-context learning ability Brown (2020); Kojima et al. (2022), and drastically longer context windows — from BART’s 1024 tokens to more than a million for modern LLMs Lee et al. (2024). As a result, recent RAG research has shifted to using LLMs Gao et al. (2023); Fan et al. (2024); in this work, we follow this trend to focus on LLM-based RAG.
@@
Dialogue generation Moghe et al. (2018); Dinan et al. (2019); Li et al. (2024) is one of the major applications of RAG, where the target response is conditioned on retrieved knowledge, and a knowledge selection step is often added to further refine the retrieved knowledge Thulke et al. (2021); Sun et al. (2023); Zhang et al. (2023). For example, Kim et al. (2020) train a knowledge selector by leveraging response information, Li et al. (2022) select knowledge from document semantic graphs, and Zhang et al. (2023) propose multi-task learning for knowledge selection and response generation. However, while these works demonstrate the advantages of their approaches through ablation studies, it is unclear whether their performance improvement via knowledge selection specifically is transferable to other datasets, domains, or tasks. Moreover, despite the popularity of LLMs for RAG, there are dramatically fewer LLM-based works that include the knowledge selection step, and it is likewise omitted in other RAG tasks Gao et al. (2023); Fan et al. (2024).
To our knowledge, there is no prior work explaining this gap in knowledge selection. The closest works to ours are Cuconasu et al. (2024) and Wu et al. (2024), which study the impact of knowledge retrieval on RAG. We hypothesize that there is a selection bias effect where knowledge selection modules may have been experimented with, but ultimately not reported, in tasks or settings where it is not helpful. In this paper, we study the general effect of knowledge retrieval and selection on LLM-based RAG via hundreds of simulations, rather than being restricted to a few specific, anecdotal retrievers or selector implementations like prior works.
@@ 3.1 Task Formulation
RAG involves three steps: (1) Knowledge retrieval, where a retriever module retrieves a set of candidate knowledge $K$ based on a query $q$. This step aims to retrieve as much knowledge that is relevant to the query as possible, balancing knowledge recall and precision. (2) Optionally, knowledge selection removes retrieved knowledge that is less relevant to further improve knowledge precision, producing $K^{\prime}\subseteq K$. (3) Finally, the generator takes the query $q$ and the selected, retrieved knowledge $K^{\prime}$ to generate the output text $r$.
The knowledge in $K$ and $K^{\prime}$ can be heterogeneous and come from multiple sources, such as external documents, knowledge graphs, or conversation histories. This RAG framework can be applied to various tasks, such as dialogue generation, question answering, fact verification, and code generation Gao et al. (2023); Fan et al. (2024). In our experiments on dialogue generation and question answering, we simulate steps 1 and 2 by creating controlled knowledge sets $K^{\prime}$ and observing the resulting generation performance of step 3.
@@ 5 Meta-Experimental Results
All of our results show that the overall RAG performance depends on the generator model: the cross-model comparisons in Figures 2 & 4 show that for stronger generator models (as measured by rank in LLM leaderboards [^1] https://huggingface.co/spaces/lmsys/chatbot-arena-leaderboard), performance across all knowledge settings are higher (Tables 1 & 2). While this result is unsurprising, we also find that the gap between the “full knowledge” and “gold knowledge” settings becomes narrower, suggesting that stronger generators are more robust to noisy input knowledge and rely less on the knowledge selector. In contrast, when the generator is weaker, any reasonable knowledge selector becomes beneficial, likely because a weak generator cannot handle noisy input and requires a selector to filter out distractor knowledge; knowledge F1 best correlates with answer F1 for Mistral-7B-Instruct in Figure 3.
@@ 6.2 Recommendations to Practitioners
Only when the knowledge recall score is high can a knowledge selector, which increases knowledge precision, be potentially helpful [^2] In a length-constrained scenario, if we consider $k$ randomly sampled knowledge sentences as a baseline, rather than the “full knowledge” setting, then the knowledge selector is more likely to be helpful due to weaker baseline performance.. Moreover, if the recall is too low, the downstream generation performance may not see any benefits from a knowledge selector with middling performance — it may even be harmed; only when the knowledge selector’s precision is very high will the overall performance improve.
@@ A.1.2 Models
We do not perform any fine-tuning or hyperparameter tuning. We set the LLMs’ temperature to 0 and ensure that our zero-shot generation prompts (Appendix A.2) are shorter than their maximum input lengths. We use the API services of OpenAI gpt-4o-mini-2024-07-18333https://platform.openai.com/docs/models/gpt-4o-mini, Together AI [^4] https://docs.together.ai/docs/chat-models meta-llama/Meta-Llama-3.1-8B-Instruct-Turbo, and mistralai/Mistral-7B-Instruct-v0.1. We report results on the first 500 examples in the HotpotQA training set and the first 100 conversations (452 wizard utterances) from the “test seen” set of WoW. Our set of experiments cost about 50 USD from OpenAI and about 50 USD from Together.ai.
@@ A.6 Length-Constrained Knowledge Selection and Answer Generation
To study the effect of length-constrained knowledge inputs on answer generation performance, we limit the number of knowledge sentences to $k$ by randomly sub-sampling $k$ sentences when more than $k$ sentences are selected by the simulated knowledge selector [^5] This random sampling does not break the experiment because our knowledge retriever and selector are simulated via random sampling, which is not the case for actual model-based retrievers and selectors.. We choose $k=3$ for both WoW and HotpotQA because their average number of candidate knowledge sentences before random sampling are 7.1 and 9.5 respectively.
@@
{endfor %}
{endif %}
{if context %}
@@
{for sentence in evidence.sentences %}
@@
{endfor %}
{endfor %}
{endif %}
@@
{for sentence in evidence.sentences %}
@@
{endfor %}
{endfor %}
{endif %}

### v4

@@ Abstract
Retrieval-augmented generation (RAG) is a powerful method for enhancing natural language generation by integrating external knowledge into a model’s output. While prior work has demonstrated the importance of improving knowledge retrieval for boosting generation quality, the role of knowledge selection, a.k.a. reranking or filtering, remains less clear. This paper empirically analyzes how knowledge selection influences downstream generation performance in RAG systems. By simulating different retrieval and selection conditions through a controlled mixture of gold and distractor knowledge, we assess the impact of these factors on generation outcomes. Our findings indicate that the downstream generator model’s capability, as well as the complexity of the task and dataset, significantly influence the impact of knowledge selection on the overall RAG system performance. In typical scenarios, improving the knowledge recall score is key to enhancing generation outcomes, with the knowledge selector providing limited benefit when a strong generator model is used on clear, well-defined tasks. For weaker generator models or more ambiguous tasks and datasets, the knowledge F1 score becomes a critical factor, and the knowledge selector plays a more prominent role in improving overall performance.
@@ 1 Introduction
Retrieval-augmented generation (RAG) has become a pivotal technique in natural language generation, enhancing a language model’s ability to produce relevant, informed output by incorporating external knowledge Gao et al. (2023); Fan et al. (2024); Gan et al. (2025). This approach complements the model’s internal knowledge, which is inherently constrained by the information that was available during training, by retrieving up-to-date, relevant information to support and inform its output.
The quality of the retrieved knowledge directly influences the quality of RAG outputs. Previous studies have consistently shown that improving knowledge retrieval leads to a direct enhancement in generation performance Dinan et al. (2019); Li et al. (2022, 2024); Wang et al. (2024); Wu et al. (2024). Similarly, effective knowledge selection, a.k.a. reranking or filtering, has been observed to improve generation quality by filtering out retrieved information that is less relevant to the generation target Kim et al. (2020); Thulke et al. (2021); Li et al. (2022); Sun et al. (2023); Zhang et al. (2023); Wang et al. (2023); Zheng et al. (2024); Wang et al. (2024); Zhao et al. (2025). However, we observe that knowledge selection is barely used for tasks other than dialogue generation, and there are notably fewer LLM-based RAG works that use knowledge selection, compared to fine-tuned RAG models (see Section 2).
@@
Therefore, in this work, we perform a systematic empirical analysis***https://github.com/jacklxc/KnowledgeSelectionSimulation of how knowledge selectors with various performances affect downstream RAG performance. By blending gold knowledge with distractor knowledge in varying ratios, we simulate different selection outcomes and examine their impact on the overall performance of RAG. We find that the generator model’s capability, as well as the complexity of the task and dataset, significantly influence RAG system performance. In typical scenarios, improving knowledge recall via the knowledge retriever is key to enhancing generation outcomes, with the knowledge selector providing limited additional benefit when a strong generator model is used on clear, well-defined tasks. For weaker generator models or more ambiguous tasks and datasets, knowledge F1 becomes a critical factor, and the knowledge selector plays a more prominent role in improving overall performance.
@@ 2 Related Work
has been extensively studied in recent years. Early works Guu et al. (2020); Lewis et al. (2020b); Shuster et al. (2021) jointly fine-tuned a dense retriever (e.g. DPR Karpukhin et al. (2020)) and generator (e.g. BART Lewis et al. (2020a)), which required dedicated training datasets. More recently, the introduction of large language models (LLMs) has made RAG more convenient to implement due to their strong generation performance, in-context learning ability Brown (2020); Kojima et al. (2022), and drastically longer context windows — from BART’s 1024 tokens to more than a million for modern LLMs Lee et al. (2024). As a result, recent RAG research has shifted to using LLMs Gao et al. (2023); Fan et al. (2024); Gan et al. (2025); in this work, we follow this trend to focus on LLM-based RAG.
@@
Dialogue generation Moghe et al. (2018); Dinan et al. (2019); Li et al. (2024) is one of the major applications of RAG, where the target response is conditioned on retrieved knowledge, and a knowledge selection step is often added to further refine the retrieved knowledge Thulke et al. (2021); Sun et al. (2023); Zhang et al. (2023). For example, Kim et al. (2020) train a knowledge selector by leveraging response information, Li et al. (2022) select knowledge from document semantic graphs, Zhang et al. (2023) propose multi-task learning for knowledge selection and response generation, and Zhao et al. (2025) proposes a multi-step reranking process for Natural Question Kwiatkowski et al. (2019) and Trivia QA Joshi et al. (2017). However, while these works demonstrate the advantages of their approaches through ablation studies, it is unclear whether their performance improvement via knowledge selection specifically is transferable to other datasets, domains, or tasks. Moreover, despite the popularity of LLMs for RAG, there are dramatically fewer LLM-based works that include the knowledge selection step, and it is likewise omitted in other RAG tasks Gao et al. (2023); Fan et al. (2024); Gan et al. (2025).
To our knowledge, there is no prior work explaining this gap in knowledge selection. The closest works to ours are Cuconasu et al. (2024), Wu et al. (2024) and Jin et al. (2025), which study the impact of knowledge retrieval on RAG. We hypothesize that there is a selection bias effect where knowledge selection modules may have been experimented with, but ultimately not reported, in tasks or settings where it is not helpful. In this paper, we study the general effect of knowledge retrieval and selection on LLM-based RAG via hundreds of simulations, rather than being restricted to a few specific, anecdotal retrievers or selector implementations like prior works.
@@ 3.1 Task Formulation
RAG involves three steps: (1) Knowledge retrieval, where a retriever module retrieves a set of candidate knowledge $K$ based on a query $q$. This step aims to retrieve as much knowledge that is relevant to the query as possible, balancing knowledge recall and precision. (2) Optionally, knowledge selection, a.k.a. reranking or filtering, removes retrieved knowledge that is less relevant to further improve knowledge precision, producing $K^{\prime}\subseteq K$. (3) Finally, the generator takes the query $q$ and the selected, retrieved knowledge $K^{\prime}$ to generate the output text $r$.
The knowledge in $K$ and $K^{\prime}$ can be heterogeneous and come from multiple sources, such as external documents, knowledge graphs, or conversation histories. This RAG framework can be applied to various tasks, such as dialogue generation, question answering, fact verification, and code generation Gao et al. (2023); Fan et al. (2024); Gan et al. (2025). In our experiments on dialogue generation and question answering, we simulate steps 1 and 2 by creating controlled knowledge sets $K^{\prime}$ and observing the resulting generation performance of step 3.
@@ 5 Meta-Experimental Results
All of our results show that the overall RAG performance depends on the generator model: the cross-model comparisons in Figures 2 & 4 show that for stronger generator models (as measured by rank in LLM leaderboards***https://huggingface.co/spaces/lmsys/chatbot-arena-leaderboard), performance across all knowledge settings are higher (Tables 1 & 2). While this result is unsurprising, we also find that the gap between the “full knowledge” and “gold knowledge” settings becomes narrower, suggesting that stronger generators are more robust to noisy input knowledge and rely less on the knowledge selector. In contrast, when the generator is weaker, any reasonable knowledge selector becomes beneficial, likely because a weak generator cannot handle noisy input and requires a selector to filter out distractor knowledge; knowledge F1 best correlates with answer F1 for Mistral-7B-Instruct in Figure 3.
@@ 6.2 Recommendations to Practitioners
Only when the knowledge recall score is high can a knowledge selector, which increases knowledge precision, be potentially helpful***In a length-constrained scenario, if we consider $k$ randomly sampled knowledge sentences as a baseline, rather than the “full knowledge” setting, then the knowledge selector is more likely to be helpful due to weaker baseline performance.. Moreover, if the recall is too low, the downstream generation performance may not see any benefits from a knowledge selector with middling performance — it may even be harmed; only when the knowledge selector’s precision is very high will the overall performance improve.
@@ A.1.1 Knowledge Sampling
We maintain the original order of the knowledge sentences from the documents provided in the datasets and do not observe a strong influence from the position of the gold knowledge sentences.
@@ A.1.2 Models
We do not perform any fine-tuning or hyperparameter tuning. We set the LLMs’ temperature to 0 and ensure that our zero-shot generation prompts (Appendix A.2) are shorter than their maximum input lengths. We use the API services of OpenAI gpt-4o-mini-2024-07-18***https://platform.openai.com/docs/models/gpt-4o-mini, Together AI***https://docs.together.ai/docs/chat-models meta-llama/Meta-Llama-3.1-8B-Instruct-Turbo, and mistralai/Mistral-7B-Instruct-v0.1. We report results on the first 500 examples in the HotpotQA training set and the first 100 conversations (452 wizard utterances) from the “test seen” set of WoW. Our set of experiments cost about 50 USD from OpenAI and about 50 USD from Together.ai.
@@ A.6 Length-Constrained Knowledge Selection and Answer Generation
To study the effect of length-constrained knowledge inputs on answer generation performance, we limit the number of knowledge sentences to $k$ by randomly sub-sampling $k$ sentences when more than $k$ sentences are selected by the simulated knowledge selector***This random sampling does not break the experiment because our knowledge retriever and selector are simulated via random sampling, which is not the case for actual model-based retrievers and selectors.. We choose $k=3$ for both WoW and HotpotQA because their average number of candidate knowledge sentences before random sampling are 7.1 and 9.5 respectively.
@@
{% endfor %}
{% endif %}
{% if context %}
@@
{% for sentence in evidence.sentences %}
@@
{% endfor %}
{% endfor %}
{% endif %}
@@
{% for sentence in evidence.sentences %}
@@
{% endfor %}
{% endfor %}
{% endif %}